OpenStudy (anonymous):

In Lake Kerr, a black bass must be less than or equal to 15 inches, or longer than or equal to 24 inches for the fish to be considered a keeper. Choose the combined inequality that describes this scenario.

OpenStudy (anonymous):

\[15 \le x \le24 \] \[x \le15 or x \ge 24 \] \[x \le24 \] none of the above

OpenStudy (anonymous):

Less than or equal to 15: \[\le 15\] or longer than or equal to 24: \[\ge 24\] \[x \le 15\ or\ x\ \ge 24\]

OpenStudy (anonymous):

is its B

OpenStudy (anonymous):

yes